OpenStudy (anonymous):

which is equivalent to 3x/x-7 - x+7/x a) 2x^2+49/x(x-7) B)x^2+49/x(x-7) c)2(x+7)/x(x-7) d)2(x-7)/x(x-7)

OpenStudy (mertsj):

\[\frac{3x}{x-7}-\frac{x+7}{x}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

You need to write both fractions with the same denominator so:

OpenStudy (mertsj):

multiply the first fraction by x/x and

OpenStudy (mertsj):

multiply the second fraction by (x-7)/(x-7)

OpenStudy (mertsj):

\[\frac{3x}{x-7}\times\frac{x}{x}-\frac{x+7}{x}\times\frac{x-7}{x-7}\]

OpenStudy (anonymous):

so you get \[\frac{ 3x ^{2} }{x ^{2-7} }\times \frac{ x ^{2-49} }{x ^{2-7} }\] right?

OpenStudy (mertsj):

\[\frac{3x^2}{x(x-7)}-\frac{x^2-49}{x(x-7)}\]

OpenStudy (mertsj):

\[\frac{3x^2-x^2+49}{x(x-7)}\]

OpenStudy (mertsj):

Collect the like terms.

OpenStudy (anonymous):

ok I think I get it so the answer is "a"??